OpenStudy (anonymous):

Madeline has an annuity that pays $7,902 at the beginning of a term twice a year. If the economy grows at a rate of 3.1% twice a year, what is the value of the annuity if she received it in a lump sum now rather than over a period of 16 years? $198,170.44 $163,868.69 $201,242.08 $158,941.50

OpenStudy (anonymous):

Someone help please??

OpenStudy (phi):

I think you can solve this using \[ \text{Future Value}= \text{payment}\cdot \frac{(1+i)^n-1}{i} \] and \[ \text{Future Value}= \text{Present Value}\cdot (1+i)^n\] n is the number of periods, 2 times a year for 16 years, or n=32 economy grows at a rate of 3.1% twice a year, so i= 0.031 per 6 months the payment is $7,902

OpenStudy (phi):

Now find the future value, using the first equation. use the future value in the second equation to find the present value, which is the answer to the question.